Volume of a gas at 35℃ and 1atmospheric pressure is 3.75 litre. If volume is to be made 3 litre at same pressure , then the temperature will be

Grade:11

2 Answers

neelam
6 years ago
P1V1/T1 = P2V2/T2 from the given problem the value of P1= P2= 1atm., V1= 3.75, V2= 3 T1= 35degree Celsius, T2=? . Putting all these values in given formula we find tha value of T2= 28degree Celsius ANSWER= 28degree celsius
Sudhir Bishnoi
6 years ago
We know that from ideal gas equation, PV=nRT,
Where T is in Kelvin
Initially P=1atm, V=3.75L, T=308.15K
putting in ideal gas equation 1×3.75=nR×308.15
                                                   nR=0.01216
 
Now P=1atm, V=3L, T have to find
                                              1×3=nRT
                                                3=0.01216×T
                                               T=246.71K                   which is answer